AnimeGhoul8863:

• Create your own piecewise function with at least two functions. Explain, using complete sentences, the steps for graphing the function. Graph the function by hand or using a graphing software of your choice (remember to submit the graph).

dude:

Okay this concept is not bad Create 2 equations One can be linear and the other quadratic

AnimeGhoul8863:

hmmmmmm 2 equations hmmmmmmmm

dude:

You can use like y=7x+1 and y=x^2

AnimeGhoul8863:

okie can i just use these ones u gave me or do you prefer me making one my self e.e

dude:

You can use mine but they came out of the top of my head XD

AnimeGhoul8863:

XD thats okie so once we have the 2 equations what next e.e

dude:

Okay just checked for an easier one to graph, let’s use y=4/5x+1 and y=1/5x^2

AnimeGhoul8863:

alright e.e

dude:

Okay let’s make the linear equation be from negative infinity to -1 The parabola would go from -1 to infinity

AnimeGhoul8863:

y=4/5x+1 this = -1? on my graph it equal 1 e.e like (0,1)

AnimeGhoul8863:

dude:

Yep So piecewise defined means that they are limited at certain points https://www.desmos.com/calculator/aisvl8r4fr So in here I made it so that the line goes until -1 and the parabola starts at -1

AnimeGhoul8863:

Ok what next

dude:

I already did the last step, but we have to describe how to draw it Do you know how to describe it? or do you need help?

AnimeGhoul8863:

it says i can use the picture of the graph thats what imma do is just screen shot the graph and add it

AnimeGhoul8863:

oh are you talking about the the steps for graphing the function.

dude:

Yeah

AnimeGhoul8863:

ohh okie then i might need some help but ill do my best to do it my self

dude:

Okay

AnimeGhoul8863:

In this graph the equations \[y=\frac{ 4}{ 5 }x+1\left\{ x \le -1 \right\}\] and \[y=\frac{ 1 }{ 5 }x^2\left\{ x \ge -1 \right\}\] are set to make the line go until -1 and then the parabola will start at -1 this would create a curve going up then line that decreases down

AnimeGhoul8863:

hows this im probably wrong but what u think

dude:

Well technically its not bad but I think teachers are looking more at the order at which they are drawn Tip with LaTeX: If you don't want it to create new large lines, use `\(\)`, not `\[\]` An example ¶, might want to change this We start out at (-1,\(\large \frac15\)) and start going left 4 and down 5 units, plot a point at (-6,- 3.8). We draw an arrow because it goes on forever. Now for the second equation we start out at (-1,\(\large \frac15\)) and substitute values into the equation like x= 0 and x=1. After that we continue drawing the parabolic shape forever since it goes on forever Note, we can limit this all from \(\mathbb D\): -5 \(\le x\le 1\) too make the response easier to explain

AnimeGhoul8863:

what does that mean "we can limit this all from D"

woolyfrog:

help me dude I don't understand

dude:

https://www.desmos.com/calculator/zv95cu5rsv We can limit how far it extends (D is the domain)

woolyfrog:

help

woolyfrog:

I only have three hours and I'm also doing other homework

AnimeGhoul8863:

We start our graph at (-1,15) when doing this it start going left 4 and down 5 units. Once we have plotted a point at (-6,- 3.8). We drew an arrow because it goes on forever. For the second equation we start out at (-1,15) and substitute values into the equation like x= 0 and x=1. After that we continue drawing the parabolic shape forever since it goes on forever. Since it goes on forever We can limit how far it extends.

AnimeGhoul8863:

@dude Hows this for the final product e.e

dude:

Good

AnimeGhoul8863:

YAY!!! THANK YEW SO MUCH DUDEEE

dude:

No problemo